OpenStudy (anonymous):

Q(n) = {4 if n = 0 Q(n) = {2 · Q(n − 1) − 3 if n > 0 Prove by induction that Q(n) = 2n + 3, for all n ≥ 0. (Induction on n.) Let f(n) = 2n + 3. Base Case: If n = 0, the recurrence relation says that Q(0) = 4, and the formula says that f(0) = 2 ? + 3 = ? , so they match. Inductive Hypothesis: Suppose as inductive hypothesis that Q(k − 1) = ? for some k > 0. nductive Step: Using the recurrence relation, Q(k) = 2 · Q(k − 1) − 3, by the second part of the recurrence relation = 2( ? ) − 3, by inductive hypothesis = (2k + 6) − 3 = ? so, by induction, Q(n) = f(n) for all n ≥ 0.

OpenStudy (zzr0ck3r):

Let f(n) = 2n + 3. Base Case: If n = 0, the recurrence relation says that Q(0) = 4, and the formula says that f(0) = 2 ? + 3 = ? , so they match. f(0) =2*0+3 = 3, I don't see how they match

OpenStudy (anonymous):

see if this makes more sense

OpenStudy (zzr0ck3r):

ahh you put 2n+3

OpenStudy (anonymous):

yeah, couldnt get the equation to look right

OpenStudy (zzr0ck3r):

the first box is 0 the second is 4 the third is 2^(k-1) + 3 the fourth is 2^(k-1)+3 the last is f(k)

OpenStudy (zzr0ck3r):

I hate to tell you the answers like this, but I have to run and I hope at least this way you will figure it out.

OpenStudy (anonymous):

Thanks it does help

OpenStudy (zzr0ck3r):

1) 0 because we want f(0) 2) IH if f(k-1) then f(k) 3) f(k-1) 4) f(k-1) 5) we are trying to show if f(k-1) then f(k) and the last step shows that

OpenStudy (anonymous):

as always, you are a huge help...I understand it now, thank you

OpenStudy (zzr0ck3r):

good deal, have a good day